# Percept, Chat, and then Adapt: Multimodal Knowledge Transfer of Foundation Models for Open-World Video Recognition

### v1

@@ 1 Introduction
Video recognition tasks are important in multimedia research, due to their wide applications in surveillance, robotics, and so on. With the rapid development of deep learning, a number of video networks have been proposed in recent yearsresnet ; r2p1d ; uniformer ; videomae ; timesformer ; pr-arclip ; st-transformer-pr ; key-frame-ar ; temporal-segment . However, these recognition models mainly work on the classical video benchmarks collected under desirable conditions. In many realistic applications, the camera-shooting environments are much more complex with low-resolution targets, bad illumination conditions, unusual video scenes, etc. In such open-world scenarios, most existing video models are not well generalized due to the lack of external domain knowledge.
@@
Alternatively, the remarkable success of large language modelsgpt3 ; llama ; minigpt4 ; chatgpt ; vicuna ; Instructgpt have driven the explosive development of various foundation modelsvideochat ; blip ; internvideo ; florence ; llava ; Text-guided ; F-SCP ; multi-embedding . Based on web-scale pre-training, these foundation models contain diversified semantic knowledge and can be adapted for low-shot generalizationflamingo ; pgc ; frozenclip ; stadapter . However, how to leverage such knowledge of these models for open-world video recognition has not been fully explored. For this reason, we propose a generic knowledge transfer paradigm, namely PCA, which can progressively exploit and integrate external knowledge of open-world videos from foundation models, by three cascade stages of Percept, Chat, Adapt. As shown in Fig.1, we treat the foundation models as a knowledge container of open-world videos, where we build PCA with the prior knowledge learned from vision, language, and multimodal models.
In the Percept stage, we first enhance poor-conditioned videos for domain gap reduction, through low-level vision models (e.g., RealBasicVSRrealbasicvsr ). Then, we extract external visual knowledge of enhanced videos, from high-level vision backbones (e.g., UniFormeruniformer ) or visual encoders of backbones (e.g. CLIPclip ). In the Chat stage, we further generate textual descriptions about predicted labels or videos, by chatting with large language models (e.g., ChatGPTchatgpt ) or visual captioners of multimodal language models (e.g., VideoChatvideochat ). Such diversified descriptions are used as external textual knowledge about videos. Finally, we introduce a distinct multimodal knowledge adaptation module in the Adapt stage, which flexibly blends various complementary knowledge into vision backbones for boosting open-world video recognition.
@@
    We evaluate PCA on three challenging tasks of open-world video recognition, e.g., low-resolution action recognition on TinyVIRATtinyvirat , dark-light action recognition on ARIDarid , and pipe anomaly detection on QV-Pipevideopipe . Extensive experiments show that our method simply achieves state-of-the-art performance on all these open-world video datasets.
@@ 2 Related work
Open World Video Recognition. In recent years, video recognition modelsslowfast ; c3d ; i3d ; ircsn ; r2p1d ; tsm ; vivit ; timesformer ; actionclip ; uniformer ; videomae have entered a period of rapid development. However, these networks are primarily designed for academic scenarios such as Kineticskinetics400 and Something-somethingssv1 datasets, rather than real-world model structures.In real-world scenarios, the captured videos contain complex environment variationsarid ; tinyvirat ; videopipe . For example, in dark environments, the low light intensity obscures important details in the video, making the action recognition task challenging. The DarkLight networkdarklight improves the accuracy of behavior recognition models in low-light conditions by incorporating a cross-attention interaction between the gamma-corrected features and the original image features. In real surveillance scenarios, when the subject is far from the camera, inevitable blurriness occurs. A weakly supervised attention mechanism is applied to highlight the regions with actiontinyvirat . QV-Pipevideopipe introduces video perception tasks to industrial applications and addresses the problem of real-world city pipeline defect recognition. In these open-world scenarios, the most common approach is the model ensemblevideopipe method with weighted addition. However, these methods can not generalize well in open-world videos with significant domain gaps.
@@
Foundation Models. Recently, language foundation models have experienced rapid development and have achieved significant improvements in text generalization capabilities. GPT-3gpt3 is a milestone model in language modeling that demonstrates the qualitative improvement brought about by increasing data and model parameters. Models such as InstructGPTInstructgpt , LLAMAllama , ChatGPTchatgpt , and Vicunavicuna make it possible for models to engage in natural conversations with humans through scaling up parameters and instruction fine-tuning. Afterwards, the innovation of language model technology has also promoted the development of visual models. The development of image foundation models also progressed rapidly. Using web-scale data and contrastive learning, image foundation modelsclip ; florence ; beit3 have achieved excellent performance in various tasks. Video foundation models such as MTVmtv , UniFormerV2uniformerv2 , and InternVideointernvideo possess exceptional video understanding capabilities. Finally, visual and textual information are combined to generate multimodal foundation modelsclip ; cosmos ; flamingo . Models such as BLIPblip ; blip2 , VideoChatvideochat , LLaVAllava , and MiniGPT-4minigpt4 have already acquired the ability to jointly process visual and natural textual information. However, how to apply the knowledge of these foundation models in open-world video understanding is still under exploration.
@@
Knowledge Adapters. Adapter modules in deep learning networks play a crucial role in various fields like knowledge transfer, large-scale pre-training, and multimodal training. By incorporating an adapter module, deep learning models have abilities to transfer pre-trained knowledge to downstream tasks and multimodal tasks. Models with adapters can be primarily categorized into two types: embedded adapter and external adapter. Embedded adapter involves modifying or adding modules within the transformer encodervit . Among them, LAVISHlavish have improved model performance by introducing lightweight multimodal fusion structures into the Transformer Encoder module. UniFormeruniformer ; uniformerv2 , ST-Adapterstadapter and BEIT-3beit3 , on the other hand, have made modifications to the self-attention and MLP modules, improving the performance of the models on multiple tasks.
@@
As for the external adapter, the original pre-trained large-scale model is preserved, and connectors are added externally for downstream tasks or other domains. CoOpCoOp incorporated additional learnable parameters as adapters to the visual branch. Clip-Adapterclipadapter , EVLfrozenclip , and Flamingoflamingo freeze the backbone network and introduce adapter modules for efficient fine-tuning. However, the models with adapter modules are not well explored in real-world video perception scenarios that involve significant domain gaps.
@@ 3.2 Percept
Figure 2: Perceptual process. We first preprocess open-world videos using low-level models to enhance the visual content and reduce the domain gap. Subsequently, we utilize visual networks to extract enhanced features and predict the confidence of all categories.
@@
For instance, in the TinyVIRATtinyvirat real low-resolution scenes, in order to obtain clearer visual information, we employed the RealBasicVSRrealbasicvsr model to perform super-resolution on the videos and enhanced certain crucial visual details. To restore some indistinguishable details in the ARID nighttime datasetarid , we brighten the video frames using gamma correction. The distinctive feature of the QV-Pipevideopipe dataset lies in its significant domain gaps. To annotate the semantic regions of the videos and reduce the domain gap, we employed the segment anythingsegmentanything model for preprocessing. By incorporating semantic segmentation masks into the original videos, we highlight some regions in the video data. Based on the aforementioned method, we obtained visual enhanced video $\widetilde{V}$. Then we utilize video foundation models $\mathscr{B}$ to extract enhanced open-world video features $F_{V}$ and predicted scores $S$. The predicted scores $S$ refer to the confidence of each category obtained after passing the visual foundation models. The whole process is formulated as
$\begin{gathered}\{F_{V},S\}=\mathscr{B}(\widetilde{V}).\end{gathered}$ (2)
It should be noted that the foundation models $\mathscr{B}$ can encompass various kinds of models, not just a single model. In order to validate the applicability of our PCA pipeline, we chose well-performing models in the field of video understanding as baselines $\mathscr{B}$. The selected models include the best-performing ones in these three real-world behavior recognition datasets, facilitating result comparison. The selection of these five models includes CNN, Transformer, and a combination of CNN and Transformer. We chose them to demonstrate that the PCA pipeline serves as a general knowledge transfer framework. The visual knowledge extractors include popular CNN models (R(2+1)Dr2p1d ), transformer models (CLIP-ViTclip , Timesformertimesformer ), as well as hybrid models (UniFormeruniformer , CLIP-3Dclip ) combining convolution and attention modules. Among them, CLIP-ViT is the visual branch of CLIPclip . CLIP-3Dclip is obtained by inserting $3\times 1\times 1$ 3D-convolutions into the transformer blocks of CLIP-ViT for better temporal information aggregation, following UniFormerV2uniformerv2 .
@@ 3.3 Chat
To save computational resources required to invoke the visual language model and obtain effective external textual knowledge, we introduced a threshold-controlled switch as shown in Fig.3. This switch allows us to choose between prompt knowledge and caption knowledge from the large visual language model, based on the confidence scores $S$ provided in the percept process. When the maximum confidence score exceeds the pre-defined threshold $\sigma$, it indicates that the augmented features yield meaningful results. In such cases, we use the prompt method $\mathscr{P}$ to obtain supplementary textual knowledge. In the TinyVIRATtinyvirat and ARIDarid datasets, we semantically expand the predicted labels. We add a subject (e.g., ”A man is”) and adverbial phrases (e.g., ”on the street”, ”in the dark”) to the category labels. If there are multiple candidate labels, we concatenate the expanded texts together. The example is shown in Fig.3. In QV-Pipevideopipe , explanatory text is published for each category. For data with multiple labels, we use ChatGPTchatgpt to integrate and summarize the explanations of these labels.
On the other hand, if none of the maximum confidence scores exceeds the threshold $\sigma$, it suggests that the percept model lacks confidence in the results obtained from visually enhanced videos. In such cases, we directly input the original video into the VideoChatvideochat $\mathscr{C}$, to obtain the external video caption $T_{c}$. Finally, we use BERTbert model $\mathscr{L}$ to extract textual features $F_{T}$ from $T_{p}$ or $T_{c}$. The process can be formulated as
@@ 4 Experiments
Datasets. We perform extensive experiments on the open-world video perception benchmarks including TinyVIRATtinyvirat , ARIDarid , and QV-Pipevideopipe . TinyVIRAT datasettinyvirat is a real-world low-resolution action recognition dataset. The resolution of the videos in the dataset ranges from 10 to 128 pixels. There are a total of 26 categories, and each video may have multiple labels. ARID datasetarid is specifically introduced for dark light scenes action recognition, consisting of 3,780 video clips categorized into 11 action classes. QV-Pipe datasetvideopipe is a specialized dataset for anomaly detection in underground water pipelines in urban areas. It is a unique dataset that exhibits significant domain differences compared to ordinary classification datasets. There are a total of 17 categories, consisting of 1 normal class and 16 defect classes.
@@
Implementation Details. Common settings for the TinyVIRAT, ARID, and QV-PIPE datasets are listed below. We evaluated the performance of the PCA pipeline using five popular models: R(2+1)Dr2p1d for CNN-style models, CLIP-ViTvit ; clip and Timesformertimesformer for transformer-style models, as well as CLIP-3D and UniFormeruniformer for CNN-transformer hybrid models. CLIP-3D is obtained by inserting 3D convolutions into ViTvit encoder module following UniFormerV2uniformerv2 . It performs well in low-resolution recognition tasks as it can encode both CLIP pretraining knowledge and temporal information. Regarding model initialization, we used publicly available pretrained weights. The R(2+1)D model used weights pretrained on the IG65mig65m dataset. The CLIP-ViT and CLIP-3D models used CLIPclip pretraining. The Timesformertimesformer and UniFormeruniformer models used Kinetics-400kinetics400 pretraining. As for the training hyperparameters, we follow most of the original settings of different modelsclip ; timesformer ; uniformer ; r2p1d . And we use AdamWadamw optimizer with momentum $\beta_{1},\beta_{2}=0.9,0.999$, and cosine decay for learning rate. All experiments were conducted on 8 NVIDIA RTX A6000 GPUs. Other different settings for each dataset are listed below.
In TinyVIRAT, the specific hyperparameters are as follows: the learning rate is set to 1e-5, the batch size is set to 72, the total number of training epochs is 90, the warmup epoch is set to 10, and the weight decay is set to 0.01. During training and testing, the images are resized to 224$\times$224 pixels using trilinear interpolation, and 8 frames are uniformly sampled. We use the F1-Score to evaluate the model performances.
@@
In ARID, most of our hyperparameters are consistent with Darklightdarklight . The learning rate is set to 1e-5, the batch size is 32, the warmup epochs is 5 and the total training epochs is 120. During training, we apply multi-scale crop and random horizontal flip augmentation. During testing, we use center-crop augmentation. The gamma correction rate is set to 1.8, and the classification head has a dropout rate of 0.5. The Top-1 accuracy and Top-5 accuracy are used for evaluation.
Table 1: Comparison with the state-of-the-art on TinyVIRATtinyvirat

### v2

# Percept, Chat, Adapt: Knowledge Transfer of Foundation Models for Open-World Video Recognition

@@ 1 Introduction
Video recognition tasks are important in multimedia research, due to their wide applications in surveillance, robotics, and so on. With the rapid development of deep learning, a number of video networks have been proposed in recent years[uniformerv2, pr-arclip, chen2022low, chen2025g, chen2025lvagent, chen2025super]. However, these recognition models mainly work on the classical video benchmarks collected under desirable conditions. In many realistic applications, the camera-shooting environments are much more complex with low-resolution targets, bad illumination conditions, unusual video scenes, etc. In such open-world scenarios, most existing video models are not well generalized due to the lack of external domain knowledge.
@@
Alternatively, the remarkable success of large language models[llama, chatgpt, vicuna, Instructgpt] have driven the explosive development of various foundation models[videochat, blip, internvideo, llava]. Based on web-scale pre-training, these foundation models contain diversified semantic knowledge and can be adapted for low-shot generalization[flamingo, chen2025top, chen2025videochat, wang2025videochat, chen2025vragent, yue2025uniflow]. However, how to leverage such knowledge of these models for open-world video recognition has not been fully explored. For this reason, we propose a generic knowledge transfer paradigm, namely PCA, which can progressively exploit and integrate external knowledge of open-world videos from foundation models, by three cascade stages of Percept, Chat, Adapt. As shown in Fig.1, we treat the foundation models as a knowledge container of open-world videos, where we build PCA with the prior knowledge learned from vision, language, and multimodal models.
In the Percept stage, we first enhance poor-conditioned videos for domain gap reduction, through low-level visual models (e.g., RealBasicVSR[realbasicvsr]). Then, we extract external visual knowledge of enhanced videos, from high-level visual models (e.g., UniFormer[uniformer]) or visual encoders of backbones (e.g. CLIP[clip]). In the Chat stage, we further generate textual descriptions about predicted labels or videos, by chatting with large language models (e.g., ChatGPT[chatgpt]) or visual captioners of multimodal language models (e.g., VideoChat[videochat]). Such diversified descriptions are used as external textual knowledge about videos. Finally, we introduce a distinct multimodal knowledge adaptation module in the Adapt stage, which flexibly blends various complementary knowledge into vision backbones for boosting open-world video recognition.
@@
    We evaluate PCA on three challenging tasks of open-world video recognition, e.g., low-resolution action recognition on TinyVIRAT[tinyvirat], dark-light action recognition on ARID[arid], and pipe anomaly detection on QV-Pipe[videopipe]. Extensive experiments show that our method simply achieves state-of-the-art performance on all these open-world video datasets.
@@ 2 Related work
Open World Video Recognition. In recent years, video recognition models[vivit, timesformer, actionclip, uniformer] have entered a period of rapid development. However, these networks are primarily designed for academic scenarios such as Kinetics[kinetics400] and Something-something[ssv1] datasets, rather than real-world model structures.In real-world scenarios, the captured videos contain complex environment variations[videopipe, tmmopenset, openworld_ar, tpami_tem_action_localization]. For example, in dark environments, the low light intensity obscures important details in the video, making the action recognition task challenging. The DarkLight network[darklight] improves the accuracy of behavior recognition models in low-light conditions by incorporating a cross-attention interaction between the gamma-corrected features and the original image features. In real surveillance scenarios, when the subject is far from the camera, inevitable blurriness occurs. A weakly supervised attention mechanism is applied to highlight the regions with action[tinyvirat]. QV-Pipe[videopipe] introduces video perception tasks to industrial applications and addresses the problem of real-world city pipeline defect recognition. In these open-world scenarios, the most common approach is the model ensemble[videopipe] method with weighted addition. However, these methods can not generalize well in open-world videos with significant domain gaps.
@@
Foundation Models. Recently, language foundation models have experienced rapid development and have achieved significant improvements in text generalization capabilities. GPT-3[gpt3] is a milestone model in language modeling that demonstrates the qualitative improvement brought about by increasing data and model parameters. Models such as InstructGPT[Instructgpt], LLAMA[llama], ChatGPT[chatgpt], and Vicuna[vicuna] make it possible for models to engage in natural conversations with humans through scaling up parameters and instruction fine-tuning. Afterwards, the innovation of language model technology has also promoted the development of visual models. The development of image foundation models also progressed rapidly. Using web-scale data and contrastive learning, image foundation models[clip, beit3] have achieved excellent performance in various tasks. Video foundation models such as UniFormerV2[uniformerv2], and InternVideo[internvideo] possess exceptional video understanding capabilities. Finally, visual and textual information are combined to generate multimodal foundation models[clip, flamingo]. Models such as BLIP[blip], VideoChat[videochat] and LLaVA[llava] have already acquired the ability to jointly process visual and natural textual information. However, how to apply the knowledge of these foundation models in open-world video understanding is still under exploration.
@@
Knowledge Adapters. Adapter modules in deep learning networks play a crucial role in various fields like knowledge transfer, large-scale pre-training, and multimodal training. By incorporating an adapter module, deep learning models have abilities to transfer pre-trained knowledge to downstream tasks and multimodal tasks. Models with adapters can be primarily categorized into two types: embedded adapter and external adapter. Embedded adapter involves modifying or adding modules within the transformer encoder[vit]. UniFormer[uniformer, uniformerv2], ST-Adapter[stadapter] and BEIT-3[beit3], on the other hand, have made modifications to the self-attention and MLP modules, improving the performance of the models on multiple tasks.
@@
As for the external adapter, the original pre-trained large-scale model is preserved, and connectors are added externally for downstream tasks or other domains. CoOp[CoOp] incorporated additional learnable parameters as adapters to the visual branch. Clip-Adapter[clipadapter], EVL[frozenclip], and Flamingo[flamingo] freeze the backbone network and introduce adapter modules for efficient fine-tuning. However, the models with adapter modules are not well explored in real-world video perception scenarios that involve significant domain gaps.
@@ 3.2 Percept
Figure 2: Percept process. We first preprocess open-world videos using low-level visual models to enhance the visual content and reduce the domain gap. Subsequently, we utilize visual networks to extract enhanced features and predict the confidence of all categories.
@@
For instance, in the TinyVIRAT[tinyvirat] real low-resolution scenes, in order to obtain clearer visual information, we employed the RealBasicVSR[realbasicvsr] model to perform super-resolution on the videos and enhanced certain crucial visual details. To restore some indistinguishable details in the ARID nighttime dataset[arid], we brighten the video frames using gamma correction. The distinctive feature of the QV-Pipe[videopipe] dataset lies in its significant domain gaps. To annotate the semantic regions of the videos and reduce the domain gap, we employed the Segment Anything[segmentanything] model for preprocessing. By incorporating semantic segmentation masks into the original videos, we highlight some regions in the video data. Based on the aforementioned method, we obtained visual enhanced video $\widetilde{V}$. Then we utilize video foundation models $\mathscr{B}$ to extract enhanced open-world video features $F_{V}$ and predicted scores $S$. The predicted scores $S$ refer to the confidence of each category obtained after passing the visual foundation models. The whole process is formulated as
[…]
It should be noted that the foundation models $\mathscr{B}$ can encompass various kinds of models, not just a single model. In order to validate the applicability of our PCA pipeline, we chose well-performing models in the field of video understanding as baselines $\mathscr{B}$. The selected models include the best-performing ones in these three real-world behavior recognition datasets, facilitating result comparison. The selection of these five models includes CNN, Transformer, and a combination of CNN and Transformer. We chose them to demonstrate that the PCA pipeline serves as a general knowledge transfer framework. The visual knowledge extractors include popular CNN models (R(2+1)D[r2p1d]), transformer models (CLIP-ViT[clip], Timesformer[timesformer]), as well as hybrid models (UniFormer[uniformer], CLIP-3D[clip]) combining convolution and attention modules. Among them, CLIP-ViT is the visual branch of CLIP[clip]. CLIP-3D[clip] is obtained by inserting $3\times 1\times 1$ 3D-convolutions into the transformer blocks of CLIP-ViT for better temporal information aggregation, following UniFormerV2[uniformerv2].
@@ 3.3 Chat
$T_{p}$ is the supplementary textual knowledge obtained from prompting. $T_{c}$ is the external video caption obtained by Videochat[videochat]. We will explain how to generate them in the following. Specifically, to save computational resources required to invoke the visual language model and obtain effective external textual knowledge, we introduced a threshold-controlled switch as shown in Fig.3. This switch allows us to choose between prompt knowledge and caption knowledge from the large visual language model, based on the confidence scores $S$ provided in the percept process. When the maximum confidence score exceeds the pre-defined threshold $\sigma$, it indicates that the augmented features yield meaningful results. In such cases, we use the prompt method $\mathscr{P}$ to obtain supplementary textual knowledge $T_{p}$. In the TinyVIRAT[tinyvirat] and ARID[arid] datasets, we semantically expand the predicted labels. We add a subject (e.g., ”A man is”) and adverbial phrases (e.g., ”on the street”, ”in the dark”) to the category labels. If there are multiple candidate labels, we concatenate the expanded texts together. The example is shown in Fig.3. In QV-Pipe[videopipe], explanatory text is published for each category. For data with multiple labels, we use ChatGPT[chatgpt] to integrate and summarize the explanations of these labels.
On the other hand, if none of the maximum confidence scores exceeds the threshold $\sigma$, it suggests that the percept model lacks confidence in the results obtained from visually enhanced videos. In such cases, we directly input the original video into the VideoChat[videochat] $\mathscr{C}$, to obtain the external video caption $T_{c}$. Finally, we use BERT[bert] model $\mathscr{L}$ to extract textual features $F_{T}$ from $T_{p}$ or $T_{c}$. $T_{p}$ is the supplementary textual knowledge obtained from the prompt method. The process can be formulated as
@@ 4 Experiments
Datasets. We perform extensive experiments on the open-world video perception benchmarks including TinyVIRAT[tinyvirat], ARID[arid], and QV-Pipe[videopipe]. TinyVIRAT dataset[tinyvirat] is a real-world low-resolution action recognition dataset. The resolution of the videos in the dataset ranges from 10 to 128 pixels. There are a total of 26 categories, and each video may have multiple labels. ARID dataset[arid] is specifically introduced for dark light scenes action recognition, consisting of 3,780 video clips categorized into 11 action classes. QV-Pipe dataset[videopipe] is a specialized dataset for anomaly detection in underground water pipelines in urban areas. It is a unique dataset that exhibits significant domain differences compared to ordinary classification datasets. There are a total of 17 categories, consisting of 1 normal class and 16 defect classes. Note that, our PCA is a generic knowledge transfer framework which aims at addressing various open-world video recognition tasks. The reason why we choose these three video benchmarks is that, these benchmarks are representative with quite different open domains in the downstream, i.e., they are suitable to verify if the proposed PCA can boost such diversified and challenging open-world video recognition, via multi-modal knowledge transfer of foundation models.
@@
Implementation Details. Common settings for the TinyVIRAT, ARID, and QV-PIPE datasets are listed below. We evaluated the performance of the PCA pipeline using five popular models: R(2+1)D[r2p1d] for CNN-style models, CLIP-ViT[vit, clip] and Timesformer[timesformer] for transformer-style models, as well as CLIP-3D and UniFormer[uniformer] for CNN-transformer hybrid models. CLIP-3D is obtained by inserting 3D convolutions into ViT[vit] encoder module following UniFormerV2[uniformerv2]. It performs well in low-resolution recognition tasks as it can encode both CLIP pretraining knowledge and temporal information. Regarding model initialization, we used publicly available pretrained weights. The R(2+1)D model used weights pretrained on the IG65m[ig65m] dataset. The CLIP-ViT and CLIP-3D models used CLIP[clip] pretraining. The Timesformer[timesformer] and UniFormer[uniformer] models used Kinetics-400[kinetics400] pretraining. As for the model hyperparameters, we follow the original settings of different models[clip, timesformer, uniformer, r2p1d]. And we use AdamW[adamw] optimizer with momentum $\beta_{1},\beta_{2}=0.9,0.999$, and cosine decay for learning rate. All experiments were conducted on 8 NVIDIA RTX A6000 GPUs. Other different settings for each dataset are listed below.
[…]
@@
In ARID, most of our hyperparameters are consistent with Darklight[darklight]. The learning rate is set to 1e-5, the batch size is 32, the warmup epochs is 5 and the total training epochs is 120. We uniformly sample 8 frames per video and apply the gamma transformation to these frames. During training, we apply multi-scale crop and random horizontal flip augmentation. During testing, we use center-crop augmentation. The gamma correction[dip] rate is set to 1.8, and the classification head has a dropout rate of 0.5. The Top-1 accuracy and Top-5 accuracy are used for evaluation.
In QV-Pipe, the learning rate, weight decay, batch size, warmup epochs, and total training epochs, are set to 1e-5, 0.2, 72, 5, and 120 respectively. Since it has large domain shifts with fine-grained anomaly categories, we used the Segment Anything (SAM) model[segmentanything] to discover detailed semantic regions in such open-world scenes. We uniformly sample 8 frames with the resolution of $\times224$. We use the segment anything model whose image encoder is MAE[mae] pretrained ViT-H[vit] as mentioned in the segment anything technical report[segmentanything]. By average fusion of semantic segmentation masks into the original video frames, we highlight the critical regions in the pipe videos for visual knowledge extraction. During training, multi-scale crop, color jitter, and random grayscale augmentation are applied. During testing, the center crop is used. The dropout rate for the classification head is 0.5. MAP (Mean Average Precision) is the evaluation metric.
4.1 Comparison to the State of the Art
Table 1: Comparison with the state-of-the-art on TinyVIRAT[tinyvirat]